A principal of $6500 is invested in an account paying an annual rate of 6%. Find the amount in the account after 6 years if the account compounded semiannually, quarterly, and monthly.
Answer by Theo(3463)

your equation is:
f = p * (1 + (i/c))^(n*c)
f = future value
p = present value
i = annual interest rate
c = number of compound intervals per year
n = number of years.

your principal is 6500.
your annual interest rate is 6%
you want to find the amount in the account after 6 years.
money is compounded semi-annually, quarterly, and monthly.

money is compounded semi-annually:
p = 6500
i = .06 (annual interest rate percent divided by 100%).
n = 6
c = 2
i/c = .06/2 = .03
n*c = 6*2 = 12
formula becomes:
f = 6500 * (1.03)^12 = 9267.45

money is compounded quarterly:
p = 6500
i = .06 (annual interest rate percent divided by 100%).
n = 6
c = 4
i/c = .06/4 = .015
n*c = 6*4 = 24
formula becomes:
f = 6500 * (1.015)^24 = 9291.77


money is compounded monthly:
p = 6500
i = .06 (annual interest rate percent divided by 100%).
n = 6
c = 12
i/c = .06/12 = .005
n*c = 6*12 = 72
formula becomes:
f = 6500 * (1.005)^72 = 9308.28781

the more compounding intervals per year, the more money you make.